# SlideSLAM: Sparse, Lightweight, Decentralized Metric-Semantic SLAM for Multi-Robot Navigation

### v6

@@ Abstract
Index Terms:
Metric-Semantic SLAM; Multi-Robot Systems; Aerial Systems: Perception and Autonomy; SLAM
@@ I Introduction
Figure 2: Metric-semantic SLAM results from seven data sequences collected by heterogeneous robots. Trajectories in different colors correspond to different data sequences. Fig. 2(a) shows a 3D reconstruction of the Pennovation campus at the University of Pennsylvania. Outdoor objects, such as vehicles, tree trunks, and light poles are mapped as shown in 2(e). Indoor objects, such as chairs, tables, and monitors are mapped as shown in Fig. 2(h). Fig. 2(b) shows the same metric-semantic map overlayed on top of an accumulated point cloud constructed by our Falcon 4 UAV. Fig. 2(c) shows an orthophoto depicting the merged metric-semantic map of three parking lots and two buildings constructed by seven robots. Fig. 2(g) and Fig. 2(h) show a zoomed-in view of one of the lab buildings, and illustrate the metric-semantic qualitative mapping performance in the indoor environments.
Table I: Related work. This table compares SlideSLAM against several state-of-the-art methods, focusing on key attributes for enabling semantics-in-the-loop autonomous exploration with robot teams. In this table, decentralized methods support operation with asynchronous observation updates from other robots. Semantic localization means that objects are used for localization between consecutive key poses. Semantic loop closure indicates that objects are explicitly used for place recognition and loop closures. Optimization over object models involves explicitly modeling objects as geometric shapes and optimizing them jointly with robot poses over time. Real-time with autonomy means that the metric-semantic SLAM system operates in real-time on-board the robots and is fully integrated into an autonomous exploration or navigation system. We develop the first metric-semantic SLAM algorithm that supports decentralized operation and is deployed to enable autonomous navigation and exploration using heterogeneous robot teams.
@@
While traditional SLAM approaches [15, 16] offer excellent accuracy in geometric perception, including estimating robot poses and reconstructing 3D geometric structures (points, surfaces, voxels), they are often insufficient to support large-scale missions, particularly when used with multiple heterogeneous robots for real-time autonomy. The challenges include managing the excessive computational load, the high storage demand for large-scale maps, handling loop closure and map merging operations. Moreover, they offer limited generalizability across different robot platforms or sensing modalities and lack semantically meaningful map representations, which are required for executing semantically meaningful tasks.
Motivated by this, recent advances in SLAM have pushed toward constructing metric-semantic maps [11, 9, 8, 17, 18]. However, they are typically limited to single-robot setups. For scalability and efficiency, an ideal system should support deployment across heterogeneous robot teams, enabling effective collaboration without relying on centralized communication or global localization infrastructure. Extending single-robot metric-semantic SLAM to decentralized operation across multiple heterogeneous robots introduces several unique challenges. These include efficient place recognition and map merging among robots, maintaining consistent and probabilistically sound sensor measurement fusion under intermittent communication, and enabling collaboration among heterogeneous platforms.
@@
Figure 3: System Diagram. Our system takes in data streams from each robot’s onboard sensors, which can be either an RGBD camera or a LiDAR, and performs instance segmentation to extract semantic object features. Meanwhile, low-level odometry, either Visual-Inertial Odometry (VIO) [19] or LiDAR-Inertial Odometry (LIO) [20], provides relative-motion estimates between consecutive key poses. Next, the metric-semantic SLAM algorithm takes in such semantic observations and relative motion estimates, and constructs a factor graph consisting of both robot pose nodes and object landmark nodes. Meanwhile, our multi-robot communication module (see Fig. 4) opportunistically leverages connectivity to share lightweight semantic observations among robots in a decentralized way. Based on this shared information, our metric-semantic place recognition algorithm constantly checks for possible inter-robot loop closures at a fixed rate. Once a loop closure is detected, the resulting transformation between each pair of robots is used to transform all observations into each robot’s reference frame. These observations are then added to their own factor graphs, forming a merged metric-semantic map. Note that the entire perception-action loop runs in a decentralized manner onboard each robot. Besides the obvious differences in control algorithms, the planning modules and the front-end processing algorithms are also different across each robot platform. This is due to the need to accommodate the differences in sensing modalities (RGBD and LiDAR), operating environments (indoor, urban, and forest), and traversal modes (ground and aerial). However, the core metric-semantic SLAM framework remains the same.
@@ II-A Metric-semantic SLAM
Unlike traditional SLAM, metric-semantic SLAM constructs a map that encodes geometric features and semantic information on objects of interest. Metric-semantic SLAM has gained significant success and popularity in the past decade [21], driven by the rapid development of deep learning techniques that extract semantic information from sensor data. A variety of map representations are used in existing metric-semantic SLAM literature. Some use dense semantic maps, such as meshes [17], volumetric maps [22, 23], surfels [24], and 2.5D grid maps [25, 26, 27]. Others use sparse object-level maps with prior information of object shapes, such as centroids [11], cubes [8], ellipsoids [9], cylinders [28], structured object models with prior shape constraints [29, 7], and mesh-based object models [30, 12]. Following the recent trends in neural implicit representation, some prior works [31, 32, 33] have also incorporated it into a map representation capable for metric-semantic SLAM, where objects are modeled implicitly using latent features from neural networks and used as localization constraints and object association. While dense metric-semantic maps are suitable for obstacle avoidance and facilitate more fine-grained modeling of objects, their computational and memory demands, especially in SWaP-constrained platforms, can be significant. Thus, sparse object-level maps are desirable for real-time downstream tasks such as active information gathering, object manipulation and multi-robot collaboration. This is because semantic information is necessary for robots to perceive their environment and execute tasks in a semantically meaningful manner. Sparsity helps robots reduce resource demands on computation, storage, and communication. A recent work in this space is ConceptGraphs [18], which builds open-vocabulary, semantically rich 3D object maps. However, ConceptGraphs is not designed for real-time, onboard operations on resource-constrained robots. Its detection and 3D object mapping pipelines require extensive computational resources. By contrast, our work seeks to develop a metric-semantic SLAM framework that utilizes sparse explicitly modeled objects and supports real-time decentralized operations within a heterogeneous robot team.
@@ II-C Multi-robot SLAM
However, the prior works mentioned above only contain geometric information during mapping. Recent efforts extend the field to metric-semantic mapping. [5] presents a system producing dense 3D semantic mesh, incorporating distributed loop closure detection and distributed robust pose-graph optimization based on Riemannian block coordinate descent [51]. However, semantic information is used primarily in the map representation rather than for localization and loop closure. Another related work, [52], proposes a distributed Pose Graph Optimization (PGO) framework based on a distributed Gaussian-Seidel approach [53] by only considering overlapping constraints among robots in PGO, and objects are extracted as landmarks. Object-level semantic landmarks, because they are viewpoint invariant, can facilitate collaboration among heterogeneous robots. Few works exist to solve the problem of collaborative SLAM using heterogeneous sensors. Among them, [54] fuses local pose graphs from different SLAM algorithms into a global pose graph. However, it is limited to vision-based sensors, which require sharing images associated with each pose for map merging. Our work proposes a more generic and efficient system towards decentralized metric-semantic SLAM that supports real-time operation on SWaP-constrained heterogeneous robot platforms with either RGBD or LiDAR sensors.
@@ II-D Semantics-in-the-loop navigation and exploration
Most autonomous navigation systems rely on geometric maps. Although such maps are reliable for obstacle avoidance, they lack semantic information that robots can leverage to improve their state estimation and planning while navigating the environment. Object-based landmarks help minimize odometry drifts, thus benefiting autonomous navigation if executed onboard in real-time [6]. Recent work goes beyond improving odometry but uses semantic information for active uncertainty reduction planning [55, 56, 3], and multi-robot collaborative planning [27]. The use of semantic information in the active SLAM is relatively uncommon [57]. Some representative work in this field includes [58, 59]. However, there has not yet been a real-time decentralized metric-semantic SLAM framework that enables a team of heterogeneous robots to explore GPS-denied environments while reasoning about the semantic information and leveraging it to facilitate multi-robot collaboration. Our work seeks to develop such a framework.
@@ III-B Dec-Metric-Semantic SLAM for multi-robot exploration
In this section, we state the general active decentralized metric-semantic SLAM problem (Dec-Metric-Semantic SLAM). Let $\mathcal{I}_{t}^{(k)}$ be all past observations that robot $k$ has accrued by time $t$, both from its own measurements, as well as from messages it has received from other robots. At time $t$, every robot $k$ executes an exploration or navigation policy:
@@
We assume $(\mathbf{w}_{t}^{(k)})_{t\leq T,k\leq K}$ and $(\mathbf{v}_{t}^{(k)})_{t\leq T,k\leq K}$ are independent random variables with zero mean and covariance $\Lambda_{k}$ and $\Gamma_{k}$, respectively. Currently, both variables $\mathbf{w}_{t}^{(k)}$ and $\mathbf{v}_{t}^{(k)}$ are set to constant values since key poses $\mathbf{x}_{t}^{(k)}$ are sampled at a fixed robot travel distance interval. However, in our implementation, we provide interface to allow scaling $\mathbf{w}_{t}^{(k)}$ according to travel distance and modify $\mathbf{v}_{t}^{(k)}$ accordingly to account for sensor and object detection noise.
@@
where $[K]:=\{1,\ldots,K\}$. For any distinct pair of robots $(i,j)$, we denote their last meeting time up to time $t$ as
@@
In principle, robot $j$ can construct ^$\hat{\mathcal{M}}_{t}^{(i)}$ from $\mathcal{I}_{t}^{(i)}$. Given the fact that the compressed representation ^$\hat{\mathcal{M}}_{t}^{(i)}$ we use is memory efficient, we directly share this information to avoid the extra computation brought about by the reconstruction step. The map ^$\hat{\mathcal{M}}_{t}^{(i)}$ is shared only for checking inter-robot loop closures and estimating the transformations. Once a valid inter-robot loop closure is established, $\mathcal{I}_{t}^{(i)}$ is used to form the updated $\mathcal{I}_{t}^{(j)}$ as mentioned in Eq. 7. $\mathcal{I}_{t}^{(j)}$ will then be used to update ^$\hat{\mathcal{M}}_{t}^{(j)}$.
@@ IV-A Approach overview
We next turn to an overview of the approach for computing the information projection operator $\Pi$, which is one of the core contributions of this paper. Unlike traditional geometric-only methods such as point cloud registration, our method leverages both geometric and semantic information to find common components of maps estimated by different robots, and then refines such estimates using both its own measurements and measurements made by other robots. We illustrate this idea with a pair of robots $k$ and $j$ at time $t$, though the argument can be readily generalized to simultaneous interactions of a larger number of robots. Robot $k$ solves the maximum likelihood state estimation problem at time $t$ as follows. For every robot $j\in\mathcal{N}_{t}^{(k)}$, we consider the message $r_{j\rightarrow k}(t)=(\mathcal{I}_{t}^{(j)},\hat{\mathcal{M}}_{t}^{(j)})$. First we determine common map observations made by robots $k$ and $j$ up to time $t$ by performing inter-robot loop-closure detection on ^$\hat{\mathcal{M}}_{t}^{(j)}$ and ^$\hat{\mathcal{M}}_{t}^{(k)}$. The output of the loop-closure module yields an estimate of the relative transformation ${}^{(k)}T_{(j)}=({}^{(k)}R_{(j)},{}^{(k)}\mathbf{t}_{(j)})\in\mathbb{SE}(3)$ between reference frames of robot $j$ and robot $k$. In particular, a point with coordinates ${}^{(j)}\textbf{p}$ in the reference frame of robot $j$ has coordinates ${}^{(k)}\textbf{p}={}^{(k)}R_{(j)}{}^{(j)}\textbf{p}+{}^{(k)}\mathbf{t}_{(j)}$ in the reference frame of robot $k$. This transformation is then used to determine which landmarks in the map of robot $j$ correspond to landmarks that also exist in the map of robot $k$. Thereafter, the trajectories of both robot $j$ and $k$ are optimized together with the union of object models in the map, taking care to associate measurements of the same landmark taken by two robots with the same variable in the factor graph:
@@ IV-B Map representation
Figure 5: Semantics-in-the-loop autonomy. The Falcon 250 UAV autonomously explores the environment while constructing a voxel map utilized for low-level navigation and a metric-semantic map used for active semantic loop closure in real time. The long-horizon exploration trajectory (thin green line) is shown alongside the dynamically feasible trajectory (thick blue curve) planned toward the next exploration waypoint.
@@ IV-C Object detection and modeling
Once a robust object instance is obtained, depending on the type of object, a shape model is fitted to the instance. Objects such as vehicles use a cuboid model, tree trunks and light poles use a cylinder model and irregularly shaped objects that cannot be properly categorized into either a cuboid or cylinder use an ellipsoid model. Having determined an appropriate model for the object instance, these models are converted into factors to be added to the factor graph for the SLAM back-end optimization.

### v7

@@ Abstract
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This is the author’s accepted version of an article in IEEE Transactions on Robotics (T-RO).
@@ I Introduction
Figure 2: Metric-semantic SLAM results from seven data sequences collected by heterogeneous robots. Trajectories in different colors correspond to different data sequences. Fig. 2(a) shows a 3D reconstruction of the Pennovation campus at the University of Pennsylvania. Outdoor objects, such as vehicles, tree trunks, and light poles are mapped as shown in 2(e). Indoor objects, such as chairs, tables, and monitors are mapped as shown in Fig. 2(h). Fig. 2(b) shows the same metric-semantic map overlayed on top of an accumulated point cloud constructed by our Falcon 4 UAV. Fig. 2(c) shows an orthophoto depicting the merged metric-semantic map of three parking lots and two buildings constructed by seven robots. Fig. 2(g) and Fig. 2(h) show a zoomed-in view of one of the lab buildings
While traditional SLAM approaches [1, 2] offer excellent accuracy in geometric perception, including estimating robot poses and reconstructing 3D geometric structures (points, surfaces, voxels), they are often insufficient to support large-scale missions, particularly when used with multiple heterogeneous robots for real-time autonomy. The challenges include managing the excessive computational load, meeting the high storage demands of large-scale maps, and handling loop closure and map merging operations. Moreover, they offer limited generalizability across different robot platforms or sensing modalities and lack semantic map representations required for executing semantically meaningful tasks.
Table I: Related work. This table compares SlideSLAM against state-of-the-art methods, focusing on key attributes for enabling semantics-in-the-loop autonomy with robot teams. In this table, decentralized methods support operation with asynchronous observation updates from other robots. Semantic localization means that objects are used for localization between consecutive key poses. Semantic loop closure indicates that objects are explicitly used for place recognition and loop closures. Optimization over object models involves explicitly modeling objects as geometric shapes and optimizing them jointly with robot poses over time. Real-time with autonomy means that the metric-semantic SLAM system operates in real-time on-board the robots and is fully integrated into an autonomous exploration or navigation system.
@@
Motivated by this, recent advances in SLAM have pushed toward constructing metric-semantic maps [13, 11, 10, 17, 18]. However, they are typically limited to single-robot setups. For scalability and efficiency, an ideal system should support deployment across heterogeneous robot teams, enabling effective collaboration without relying on centralized communication or global localization infrastructure. Extending single-robot metric-semantic SLAM to decentralized operation across multiple heterogeneous robots introduces several unique challenges. These include performing efficient place recognition and map merging among robots, maintaining consistent and probabilistically sound sensor measurement fusion under intermittent communication, and enabling collaboration among heterogeneous platforms.
@@ II Related Work
Figure 3: System Diagram. Our system takes in data streams from each robot’s onboard sensors, which can be either an RGBD camera or a LiDAR, and performs instance segmentation to extract semantic object features. Meanwhile, low-level odometry, either Visual-Inertial Odometry (VIO) or LiDAR-Inertial Odometry (LIO), provides relative-motion estimates between consecutive key poses. Next, the metric-semantic SLAM algorithm takes in such semantic observations and relative motion estimates, and constructs a factor graph consisting of both robot pose nodes and object landmark nodes. Meanwhile, our multi-robot communication module (see Fig. 4) opportunistically leverages connectivity to share lightweight semantic observations among robots in a decentralized way. Based on this shared information, our metric-semantic place recognition algorithm constantly checks for possible inter-robot loop closures at a fixed rate. Once a loop closure is detected, the resulting transformation between each pair of robots is used to transform all observations into each robot’s reference frame. These observations are then added to their own factor graphs, forming a merged metric-semantic map. Note that the entire perception-action loop runs in a decentralized manner onboard each robot. Besides the obvious differences in control algorithms, the planning modules and the front-end processing algorithms are also different across each robot platform. This is due to the need to accommodate the differences in sensing modalities (RGBD and LiDAR), operating environments (indoor, urban, and forest), and traversal modes (ground and aerial). However, the core metric-semantic SLAM framework remains the same.
@@ II-A Metric-semantic SLAM
Unlike traditional SLAM, metric-semantic SLAM constructs a map that encodes geometric features and semantic information on objects of interest. Metric-semantic SLAM has gained significant success and popularity in the past decade [19], driven by the rapid development of deep learning techniques that extract semantic information from sensor data. A variety of map representations are used in existing metric-semantic SLAM literature. Some use dense semantic maps, such as meshes [17], volumetric maps [20, 21], surfels [22], and 2.5D grid maps [23, 24, 25]. Others use sparse object-level maps with prior information of object shapes, such as centroids [13], cubes [10], ellipsoids [11], cylinders [26], structured object models with prior shape constraints [27, 9], and mesh-based object models [28, 14]. Following the recent trends in neural implicit representation, some prior works [29, 30, 31] have also incorporated it into a map representation capable of metric-semantic SLAM, where objects are modeled implicitly using latent features from neural networks and used as localization constraints and for object association. While dense metric-semantic maps are suitable for obstacle avoidance and facilitate more fine-grained modeling of objects, their computational and memory demands, especially in SWaP-constrained platforms, can be significant. Thus, sparse object-level maps are desirable for real-time downstream tasks such as active information gathering, object manipulation and multi-robot collaboration. This is because semantic information is necessary for robots to perceive their environment and execute tasks in a semantically meaningful manner. Sparsity helps robots reduce resource demands on computation, storage, and communication. A recent work in this space is ConceptGraphs [18], which builds open-vocabulary, semantically rich 3D object maps. However, ConceptGraphs is not designed for real-time, onboard operations on resource-constrained robots. Its detection and 3D object mapping pipelines require extensive computational resources. By contrast, our work seeks to develop a metric-semantic SLAM framework that utilizes sparse explicitly modeled objects and supports real-time decentralized operations within a heterogeneous robot team.
@@ II-C Multi-robot SLAM
However, the prior works mentioned above only contain geometric information during mapping. Recent efforts extend the field to metric-semantic mapping. [7] presents a system producing a dense 3D semantic mesh, incorporating distributed loop closure detection and distributed robust pose-graph optimization based on Riemannian block coordinate descent [49]. However, semantic information is used primarily in the map representation rather than for localization and loop closure. Another related work, [50], proposes a distributed Pose Graph Optimization (PGO) framework based on a distributed Gaussian-Seidel approach [51] by considering only overlapping constraints among robots in the optimization and extracting objects as landmarks. Object-level semantic landmarks, because they are viewpoint invariant, can facilitate collaboration among heterogeneous robots. Few works exist to solve the problem of collaborative SLAM using heterogeneous sensors. Among them, [52] fuses local pose graphs from different SLAM algorithms into a global pose graph. However, it is limited to vision-based sensors, which require sharing images associated with each pose for map merging. Our work proposes a more generic and efficient system towards decentralized metric-semantic SLAM that supports real-time operation on SWaP-constrained heterogeneous robot platforms with either RGBD or LiDAR sensors.
@@ II-D Semantics-in-the-loop navigation and exploration
Most autonomous navigation systems rely on geometric maps. Although such maps are reliable for obstacle avoidance, they lack semantic information that robots can leverage to improve their state estimation and planning while navigating the environment. Object-based landmarks help minimize odometry drift, thus benefiting autonomous navigation if executed onboard in real-time [8]. Recent work goes beyond improving odometry but uses semantic information for active uncertainty reduction planning [53, 54, 5], and multi-robot collaborative planning [25]. The use of semantic information in active SLAM is relatively uncommon [55]. Some representative work in this field includes [56, 57]. However, there has not yet been a real-time decentralized metric-semantic SLAM framework that enables a team of heterogeneous robots to explore GPS-denied environments while reasoning about the semantic information and leveraging it to facilitate multi-robot collaboration. Our work seeks to develop such a framework.
@@ III-B Dec-Metric-Semantic SLAM for multi-robot exploration
In this section, we state the general active decentralized metric-semantic SLAM problem (Dec-Metric-Semantic SLAM). Let $\mathcal{I}_{t}^{(k)}$ be all past observations that robot $k$ has accrued by time $t$, from its own measurements, as well as from messages it has received from other robots. At time $t$, every robot $k$ executes an exploration or navigation policy:
@@
We assume $(\mathbf{w}_{t}^{(k)})_{t\leq T,k\leq K}$ and $(\mathbf{v}_{t}^{(k)})_{t\leq T,k\leq K}$ are independent random variables with zero mean and covariance $\Lambda_{k}$ and $\Gamma_{k}$, respectively. Currently, both variables $\mathbf{w}_{t}^{(k)}$ and $\mathbf{v}_{t}^{(k)}$ are set to constant values since key poses $\mathbf{x}_{t}^{(k)}$ are sampled at a fixed robot travel distance interval. However, in our implementation, we provide an interface to allow scaling $\mathbf{w}_{t}^{(k)}$ according to travel distance and modify $\mathbf{v}_{t}^{(k)}$ accordingly to account for sensor and object detection noise.
@@
where $K]:=\{1,\ldots,K\}$. For any distinct pair of robots $(i,j)$, we denote their last meeting time up to time $t$ as
@@
In principle, robot $j$ can construct $\hat{\mathcal{M}}_{t}^{(i)}$ from $\mathcal{I}_{t}^{(i)}$. Given the fact that the compressed representation $\hat{\mathcal{M}}_{t}^{(i)}$ we use is memory efficient, we directly share this information to avoid the extra computation brought about by the reconstruction step. The map $\hat{\mathcal{M}}_{t}^{(i)}$ is shared only for checking inter-robot loop closures and estimating the transformations. Once a valid inter-robot loop closure is established, $\mathcal{I}_{t}^{(i)}$ is used to form the updated $\mathcal{I}_{t}^{(j)}$ as mentioned in Eq. 7. $\mathcal{I}_{t}^{(j)}$ will then be used to update $\hat{\mathcal{M}}_{t}^{(j)}$.
@@ IV-A Approach overview
We next turn to an overview of the approach for computing the information projection operator $\Pi$, which is one of the core contributions of this paper. Unlike traditional geometric-only methods such as point cloud registration, our method leverages both geometric and semantic information to find common components of maps estimated by different robots, and then refines such estimates using both its own measurements and measurements made by other robots. We illustrate this idea with a pair of robots $k$ and $j$ at time $t$, though the argument can be readily generalized to simultaneous interactions of a larger number of robots. Robot $k$ solves the maximum likelihood state estimation problem at time $t$ as follows. For every robot $j\in\mathcal{N}_{t}^{(k)}$, we consider the message $r_{j\rightarrow k}(t)=(\mathcal{I}_{t}^{(j)},\hat{\mathcal{M}}_{t}^{(j)})$. First we determine common map observations made by robots $k$ and $j$ up to time $t$ by performing inter-robot loop-closure detection on $\hat{\mathcal{M}}_{t}^{(j)}$ and $\hat{\mathcal{M}}_{t}^{(k)}$. The output of the loop-closure module yields an estimate of the relative transformation $^{(k)}T_{(j)}=({}^{(k)}R_{(j)},{}^{(k)}\mathbf{t}_{(j)})\in\mathbb{SE}(3)$ between reference frames of robot $j$ and robot $k$. In particular, a point with coordinates $^{(j)}\textbf{p}$ in the reference frame of robot $j$ has coordinates $^{(k)}\textbf{p}={}^{(k)}R_{(j)}{}^{(j)}\textbf{p}+{}^{(k)}\mathbf{t}_{(j)}$ in the reference frame of robot $k$. This transformation is then used to determine which landmarks in the map of robot $j$ correspond to landmarks that also exist in the map of robot $k$. Thereafter, the trajectories of both robot $j$ and $k$ are optimized together with the union of object models in the map, taking care to associate measurements of the same landmark taken by two robots with the same variable in the factor graph:
@@ IV-C Object detection and modeling
Once a robust object instance is obtained, depending on the type of object, a shape model is fitted to the instance. Objects such as vehicles use a cuboid model, tree trunks and light poles use a cylinder model, and irregularly shaped objects that cannot be properly categorized into either a cuboid or cylinder use an ellipsoid model. Having determined an appropriate model for the object instance, these models are converted into factors to be added to the factor graph for the SLAM back-end optimization.